OpenStudy (anonymous):

given that f(x)=2x+1 and g(x)=-5x+2 solve for f(g(x) when x=3 A.-91 B.-33 C.-25 D.35

OpenStudy (anonymous):

so start by evaluating the inner function, in this case its g(x)=-5x+2 therefore g(3) = -5*3+2 = -15+2=-13 then plug that result into your outer function, f(x) = 2x+1 f(-13) = 2*-13+1 = -26+1=-25

OpenStudy (jdoe0001):

\(\bf f(x)=2x+1\qquad {\color{brown}{ g(x)}}=-5x+2 \\ \quad \\ f(\quad {\color{brown}{ g(x)}}\quad )=2{\color{brown}{ g(x)}}+1\to 2({\color{brown}{ -5x+2}})+1 \\ \quad \\ f(\quad {\color{brown}{ g({\color{blue}{ 3}})}}\quad )=2{\color{brown}{ g(x)}}+1\to 2({\color{brown}{ -5({\color{blue}{ 3}})+2}})+1\)

OpenStudy (anonymous):

given that f(x)=\[^{2}\]+3x+6 and g(x)=x-3 solve for f(g(x)) when x=1 ---------- 2 A.-1 B.0 C.3 4.4